The Vividh Bharati station of All India Radio, Delhi, broadcasts on a frequency of 1,368 kHz (kilo hertz). Calculate the wavelength of the electromagnetic radiation emitted by transmitter. Which part of the electromagnetic spectrum does it belong to?

Grade:12

3 Answers

Bhiniksha Mandlik
6 years ago
Delhi broadcasts at frequency(n)= 1368000HzFor wavelength of electromagnetic radiation emitted by the transmitter,Wavelength =C/V =3*10^8/1368*10^3. =0.00219 * 10^5 =219m
Arsalan
5 years ago
 
frequency of signal= 1368000Hz
wavelength of signal=c/v
wavelength=3x10^8/1368x10^3=0.00219 x10^5=219m
ankit singh
3 years ago

Answer

 

Wavelength =λ
frequency = ν = 1,368 kHZ 
     c = speed of light =v×λ
   λ=c/λ=3×108/1368×103
     = 219 meters
It belongs the far infrared region or radio wave region.